OpenStudy (olivya999):

Find the inverse of the function. f(x) = 6x3 + 3

OpenStudy (sparrow2):

is it x^3?

OpenStudy (olivya999):

f-1(x) = cube root of quantity x plus three divided by six f-1(x) = cube root of quantity x minus three divided by six Not a one-to-one function f-1(x) = cube root of x divided by six minus three

OpenStudy (olivya999):

@sparrow2

OpenStudy (sparrow2):

second

OpenStudy (jungleman03):

is ther multiple choice?

OpenStudy (olivya999):

yeah those are the multiple choices

OpenStudy (sparrow2):

it is 2

OpenStudy (olivya999):

ok thanks @sparrow2